# Galactic transient sources with the Cherenkov Telescope Array

### v1

@@ 4.2 Low-mass X-ray binaries
LMXBs harbor a low-mass companion star and a black hole (or an accreting neutron star), object tightly connected to jet launching that are responsible for the non-thermal multi-wavelength emission (see review in Chaty, 2022). Up to now, no LMXB has been detected at HE (apart from tMSPs) and only strong hints of emission at HE have been reported in V404 Cyg. No LMXB has neither been detected at VHE by any IACT. (see e.g. Aleksić et al. 2011; Ahnen et al. 2017a; H. E. S. S. Collaboration et al. 2018). The most recent X-ray outburst of a black hole LMXB (BH-LMXB) which was followed up by the IACTs MAGIC, H.E.S.S. and VERITAS was that of MAXI J1820070, without detecting any VHE emission (Abe et al., 2022). We examine here if CTA will be able to detect such a similar exceptionally bright outburst but for a hypothetical source located within a maximal distance of 4 kpc from Earth. Based on the theoretical lepto-hadronic model of Kantzas et al. (2022), used since the modeled LMXB can be considered a canonical source, we perform a number of simulations where we rescale the predicted VHE emission for a number of different jet inclination angles between 5 and . We perform each simulation for a number of different hypothetical sources at different distances within 2 and 4 kpc.
In Fig. 13, we show the predicted VHE flux for a BH-LMXB with inclination angle of assuming that the emission persists in three different energy bins between 0.1 and 10 TeV, for at least 2 weeks, and compare it to the CTA sensitivity curves (see Fig. 1).
We overall see that CTA will be able to detect an outburst similar to MAXI J1820070 in the sub-TeV regime within a few tens of minutes if the LMXB is located closer than kpc at energies TeV. The inclination angle of the LMXB we assume here is relatively small compared to the average value between and (see, e.g., Tetarenko et al., 2016), but LMXBs with an inclination angle greater than fail to be detected within the first hour of observations (Fig. 14). Sources with an inclination angle less than could be observed within a few minutes, such as the case of MAXI J1836194 or V4641 Sgr, both microblazar candidates (Russell et al., 2014; Gallo et al., 2014).
@@
We carried out a 50-h CTA simulated observation for V404 Cyg with the same setup described in Section 4.1.2: 100 GeV – 1 TeV simulated photons with a multi-source approach. The CTA input spectral model for V404 Cyg is a simple extension of the power-law spectrum observed by Fermi-LAT during the 2015-June flaring activity and reported by Loh et al. (2016) (assuming: prefactor ph , index = 3.5 and pivor energy = 1 TeV). The resulting spectrum from our simulation is shown in Fig. 15, together with the non-thermal HE spectra observed during the 2015-June flare. Thus, if we assume the same spectral trend as observed by the HE gamma-ray detectors, we expect no detection with CTA in a 50-h observation. This is in agreement with the simulations on LMXBs described in Subsection 4.2.
@@ 4.3 Transitional millisecond pulsars
Out of the three confirmed tMSPs, only PSR J1023+0038 is currently in the LMXB state, whereas XSS J12274853 and IGR J182452452 are currently in the RMSP state. As previously mentioned, other candidates were found through X-ray peculiar variability and association with Fermi-LAT sources (see review by Papitto & de Martino, 2022). Particularly interesting are the two confirmed tMSPs PSR J1023+0038 and XSS J1227-485 that when in LMXB state they were found by Fermi-LAT with a luminosity of about 1034 erg s-1 in the energy range 0.1–10 GeV, which is up to ten times brighter than the levels observed during the RMSP state (Papitto & Torres, 2015; Torres et al., 2017). This fact makes them particularly interesting for a possible detection with CTA. In this Section, we estimate the chances of detecting these two tMSPs with CTA given also their relatively close distance of about 1.5 kpc.
@@ 4.3.1 PSR J1023+0038
This tMSP was initially detected as a variable source in the radio band (Bond et al., 2002) and showing clear characteristics of an accretion disc around the compact object in the optical band. Later, Thorstensen & Armstrong (2005) suggested PSR J1023+0038 as an NS-LMXB. The observations did not reveal an accretion disc but the existence of a strong irradiation on the optical star from an unseen companion. The compact object was identified as a 1.69 millisecond radio pulsar in a 4.75 hr orbit around a 0.2 M⊙ companion star (Archibald et al., 2009b). In June 2013 the source came back to a LMXB state, where it has remained until now, and the radio pulsar signal switched off (Stappers et al., 2013). During the LMXB state, PSR J1023+0038 shows a peculiar behaviour in X-rays: it exhibits frequent modes switching between three different X-ray levels, dubbed high, low and flaring (Bogdanov et al., 2015). The HE gamma-ray emission detected by Fermi-LAT has been reported to brighten by a factor of 5 after the transition (Stappers et al., 2013). The average Fermi-LAT spectrum is described by a power-law with index 1.8 and a cutoff at an energy of 2.3 GeV according to Takata et al. (2014) and by a power-law with index 2 and an energy cutoff at 3.7 GeV, being the significance of the cutoff is 4.3 level according to Torres et al. (2017). Neither pulsations nor steady emission were found in the VHE regime (Aliu et al., 2016). To test the capability of CTA to detect emission from this source we first studied the HE gamma-ray emission from Fermi-LAT during the LMXB state (2013-2021), in order to obtain the spectral parameters of the source. The two spectral models that have been considered for the CTA simulations are a logparabola and a broken power-law. The power-law with the exponential cutoff is not further considered because the cutoff at a few GeV would prevent a detection at VHE; whereas, a simple power-law extending in the energy range from GeV up to 1 TeV appears physically difficult to achieve. This is compatible by the low significance of the results by Takata et al. (2014) and Torres et al. (2017).
For the broken power-law: prefactor = (0.060.01) 10-10 photons cm-2 s-1 MeV-1, Index1 = -2.120.03, Index2 = -2.910.06, Energy break = (1.150.09) GeV. For the logparabola: prefactor = (0.340.007) 10-10 photons cm-2 s-1 MeV-1, = 2.230.02, = 0.160.02, Energy break = 0.524 GeV. Batches of 100 simulations were run using both CTA-N and CTA-S IRFs for 50, 100 and 200 hours of observations; we binned the simulated data into initial 20 logarithmic energy bins considering an energy range of 0.03-100 TeV. The resulting spectrum of PSR J1023+0038 is shown in Fig. 16 and, for simplicity, only the results with the highest statistic are reported (200 h observations). The full analysis results for the broken power-law model are reported in Table 6. We find that long integration times are needed to detect this tMSP, with at least 100 h for CTA-N and 50 h for CTA-S.
The fit with a logparabola model did not converge.
@@ 4.3.2 XSS J1227-48538
Before the transition to the radio state, the gamma-ray emission was a factor of 2 larger (Torres et al., 2017). The Fermi-LAT analysis performed for the period in which the source was in the sub-luminous disk state (2008-2012) provides results consistent with those reported by Xing & Wang (2015) and t2017ApJ...836...68T: XSS J1227-48538 is best described by a power-law with a cutoff at Ecut=5.3 GeV (at 3.4) and a spectral index of 2.
Similarly to PSR J1023+0038, the two spectral models considered for the CTA simulations are the logparabola and the broken power-law. As input models for the CTA simulations we considered the output from the Fermi-LAT study. For the broken power-law: prefactor = (2.710.71) 10-12 photons cm-2 s-1 MeV-1, Index1 = -2.230.04, Index2 = -2.770.10, Energy break = (1.321.44) GeV. For the logparabola: prefactor = (3.260.91) 10-11 photons cm-2 s-1 MeV-1, = 2.280.05, = 0.090.02, Energy break = (4.455.42) GeV. We performed batches of 100 simulations using only the CTA-S site IRFs and 50, 100 and 200 h of observations. The simulated data were binned into 20 logarithmic energy bins in an energy range of 0.03-100 TeV.
The resulting spectrum of XSS J1227-48538 is shown in Figure 17 and, for simplicity, we reported only the results with the highest statistics. The full analysis results are reported in Table 7 . The source could be detected with CTA-S at 8.14 with 50 h of observation. As for PSR J1023+0038, simulations considering the logparabola model failed to calculate the ULs.
Our simulations prove that the detection of the spectral component seen by Fermi-LAT from of close-by tMSPs when in a disc state could be possible with long exposures, provided that the emission has no cutoff at few GeVs. While this will likely not allow to catch fully a transition if lasting less than several days, here we demonstrate that once a transition has occurred CTA could be able to detect such type of sources, identifying tMSPs as VHE emitters for the first time. Also, if there are additional components (not considered here, such as magnetic reconnection of pulsar wind) it could be possible to detect changes in the HE flux along transitions.
@@ 4.4 Flares in PWNe: the Crab Nebula
CTA will cover a fundamental energy range to understand the origin of these flares: on the one hand, the low energy threshold will allow the sampling of the Fermi-LAT spectral shape at few tens of GeV of synchrotron nature, providing important clues on the acceleration and emission processes; on the other hand, the excellent sensitivity in the TeV regime will serve to explore the IC component that might arise at a detectable level from the electron population behind the MeV flares, off-scattering soft photon fields.
@@
To evaluate the capability of CTA to detect Crab flares, we performed simulations of the SEDs both in flaring and steady (non-flaring) states of the nebula. We simulated flares of different spectral characteristics starting from a parent particle population, varying the physical properties of the environment. In particular, we simulated three types of flares: a very bright flare with similar flux (at hundreds of MeV) to the one observed by Fermi-LAT in April 2011 (Buehler et al., 2012; Striani et al., 2013) which is the flare with largest flux to date, and two dimmer flares corresponding to the first one re-scaled by a factor 0.5 and 0.1. Since no spectral variability has been reported at HE, we assume the same spectral model for these dimmer flares. The simulations of the nebula in flaring and steady states are performed for the CTA-N with the methods and tools presented in previous works (Mestre et al., 2020, 2021). The electron population was simulated with a fixed index () of 2.5, to guarantee the detection in the TeV regime (e.g., from 1.25 TeV to 50 TeV) of the brightest model of flare (see Fig. 6 of Mestre et al. 2021) in less than 10 h (see Table 8). The different flare models are computed for a magnetic field (B) in the acceleration region ranging from 100 G to 1 mG and compared (see section 2 of Mestre et al. 2021) to the steady nebula SED in both tens of GeV (e.g., from 20 GeV to 200 GeV) and TeV regimes. We obtained the simulations of the Crab nebula SED in steady state from Mestre et al. 2020. To compare the flaring and steady nebula simulations, we computed the mean total expected excess (e.g., counts from the source after background subtraction), both in flaring and steady state, in 21 bins of energy from 12 GeV to 200 TeV with observation times ranging from a few minutes to 500 h. We compared the excess distributions using the Pearson’s chi-squared test, corresponding the null hypothesis () to the steady state. Then, we consider the flare implies a detectable flux level if is rejected at 95 % confidence level (CL).
@@
The simulations performed show that the different models of flare are best detected in the GeV regime and in particular in less than an hour at energies below 200 GeV, see Fig. 18. In the TeV regime, flares dimmer than April 2011 flare by a factor 0.5 (at hundreds of MeV) would be detected in less than 10 h for B G (see Table 8). Some of the models considered imply an energy in TeV electrons larger than erg, being the synchrotron cooling time, and erg s-1, the luminosity of the nebula in gamma rays (Rudak & Dyks, 1998). However, note that the energy in electrons available in the nebula is not limited to the one computed above if particle re-acceleration takes place, which would introduce additional boosts in the electron population energy reservoir. The simulations performed, together with previous results reported (see Mestre et al., 2021), provide excellent prospects for detecting flares from the Crab Nebula with CTA, especially for the LST subarray, featuring the best sensitivity at energies of a few tens of GeV.

### v2

@@ 4.2 Low-mass X-ray binaries
LMXBs harbor a low-mass companion star and a black hole (or an accreting neutron star), object tightly connected to jet launching that are responsible for the non-thermal multi-wavelength emission (see review in Chaty, 2022). Up to now, no LMXB has been detected at HE (apart from tMSPs) and only strong hints of emission at HE have been reported in V404 Cyg. No LMXB has neither been detected at VHE by any IACT. (see e.g. Aleksić et al. 2011; Ahnen et al. 2017a; H. E. S. S. Collaboration et al. 2018). The most recent X-ray outburst of a black hole LMXB (BH-LMXB) which was followed up by the IACTs MAGIC, H.E.S.S. and VERITAS was that of MAXI J1820070, without detecting any VHE emission (Abe et al., 2022). We examine here if CTAO will be able to detect such a similar exceptionally bright outburst but for a hypothetical source located within a maximal distance of 4 kpc from Earth. Based on the theoretical lepto-hadronic model of Kantzas et al. (2022), used since the modeled LMXB can be considered a canonical source, we perform a number of simulations where we rescale the predicted VHE emission for a number of different jet inclination angles between 5 and . We perform each simulation for a number of different hypothetical sources at different distances within 2 and 4 kpc.
In Fig. 18, we show the predicted VHE flux for a BH-LMXB with inclination angle of assuming that the emission persists in three different energy bins between 0.1 and 10 TeV, for at least 2 weeks, and compare it to the CTAO sensitivity curves (see Fig. 1).
We overall see that CTAO will be able to detect an outburst similar to MAXI J1820070 in the sub-TeV regime within a few tens of minutes if the LMXB is located closer than kpc at energies TeV. The inclination angle of the LMXB we assume here is relatively small compared to the average value between and (see, e.g., Tetarenko et al., 2016), but LMXBs with an inclination angle greater than fail to be detected within the first hour of observations (Fig. 19). Sources with an inclination angle less than could be observed within a few minutes, such as the case of MAXI J1836194 or V4641 Sgr, both microblazar candidates (Russell et al., 2014; Gallo et al., 2014).
@@
We carried out a 50-h CTAO simulated observation for V404 Cyg with the same setup described in Section 4.1.2: 100 GeV – 1 TeV simulated photons with a multi-source approach. The CTAO input spectral model for V404 Cyg is a simple extension of the power-law spectrum observed by Fermi-LAT during the 2015-June flaring activity and reported by Loh et al. (2016) (assuming: prefactor ph , index = 3.5 and pivot energy = 1 TeV). The resulting spectrum from our simulation is shown in Fig. 20, together with the non-thermal HE spectra observed during the 2015-June flare. Thus, even if we assume the same spectral trend as observed by the HE gamma-ray detectors, we expect no detection with CTAO in a 50-h observation. This is in agreement with the simulations on LMXBs described in Subsection 4.2.
@@ 4.3 Transitional millisecond pulsars
Out of the three confirmed tMSPs, only PSR J1023+0038 is currently in the LMXB state, whereas XSS J12274853 and IGR J182452452 are currently in the RMSP state. As previously mentioned, other candidates were found through X-ray peculiar variability and association with Fermi-LAT sources (see review by Papitto & de Martino, 2022). Particularly interesting are the two confirmed tMSPs PSR J1023+0038 and XSS J1227-485 that when in LMXB state they were found by Fermi-LAT with a luminosity of about 1034 erg s-1 in the energy range 0.1–10 GeV, which is up to ten times brighter than the levels observed during the RMSP state (Papitto & Torres, 2015; Torres et al., 2017). This fact makes them particularly interesting for a possible detection with CTAO. In this Section, we estimate the chances of detecting these two tMSPs with CTAO given also their relatively close distance of about 1.5 kpc.
@@ 4.3.1 PSR J1023+0038
This tMSP was initially detected as a variable source in the radio band (Bond et al., 2002) and showing clear characteristics of an accretion disc around the compact object in the optical band. Later, Thorstensen & Armstrong (2005) suggested PSR J1023+0038 as an NS-LMXB. The observations did not reveal an accretion disc but the existence of a strong irradiation on the optical star from an unseen companion. The compact object was identified as a 1.69 millisecond radio pulsar in a 4.75 hr orbit around a 0.2 M⊙ companion star (Archibald et al., 2009b). In June 2013 the source came back to a LMXB state, where it has remained until now, and the radio pulsar signal switched off. During the LMXB state, PSR J1023+0038 shows a peculiar behaviour in X-rays: it exhibits frequent modes switching between three different X-ray levels, dubbed high, low and flaring (Bogdanov et al., 2015). The HE gamma-ray emission detected by Fermi-LAT has been reported to brighten by a factor of 5 after the transition. The average Fermi-LAT spectrum is described by a power-law with index 1.8 and a cutoff at an energy of 2.3 GeV according to Takata et al. (2014) and by a power-law with index 2 and an energy cutoff at 3.7 GeV, the significance of the cutoff is 4.3 level according to Torres et al. (2017). Neither pulsations nor steady emission were found in the VHE regime (Aliu et al., 2016). To test the capability of CTAO to detect emission from this source we first studied the HE gamma-ray emission from Fermi-LAT during the LMXB state (2013-2021), in order to obtain the spectral parameters of the source. For the Fermi-LAT analysis, all photons in the energy range 0.1-300 GeV included in a circular region of 10 degrees centered on the source were considered. The binned likelihood analysis was performed using 20 energy bins. The two spectral models that have been considered for the CTAO simulations are a logparabola and a broken power-law. We considered only these two models, because the simulation with the power law with an exponential cut-off model did not return any detectable spectral bin in the VHE range (the cut-off is at very low energies, a few GeV). On the other hand a simple power-law extending in the energy range from GeV up to 1 TeV appears physically difficult to achieve. This is compatible by the low significance of the results by Takata et al. (2014) and Torres et al. (2017).
The spectral input parameters inferred from the analysis of the Fermi-LAT data in the accretion phase are for the broken power-law: prefactor Pf = (0.060.01) 10-10 photons cm-2 s-1 MeV-1, = -2.120.03, = -2.910.06, Energy break = (1.150.09) GeV; while for the logparabola: prefactor Pf = (0.340.007) 10-10 photons cm-2 s-1 MeV-1, = 2.230.02, = 0.160.02, Energy break = 0.524 GeV.
Batches of 100 simulations were run using both CTAO-N and CTAO-S IRFs for 50, 100 and 200 hours of observations; we binned the simulated data into initial 20 logarithmic energy bins considering an energy range of 0.03-100 TeV.
The resulting spectrum of PSR J1023+0038 is shown in Fig. 21 and, for simplicity, only the results with the highest statistic are reported (200 h observations). The full analysis results for the broken power-law model are reported in Table 5. The parameters obtained from the unbinned analysis are:
Pf = 7.31 10-12 photons cm-2 s-1 MeV-1, = -2.12, = -2.91, Energy break = 1.15 GeV. We find that long integration times are needed to detect this tMSP, with at least 100 h for CTAO-N and 50 h for CTAO-S.
We tried to also fit the simulation obtained with a logparabola model but the fit did not converge, despite several attempts changing various parameters (prefactor, ROI, etc.). This is likely caused by the very low statistics of the simulated spectrum.
@@ 4.3.2 XSS J1227-48538
Before the transition to the radio state, the gamma-ray emission was a factor of 2 larger (Torres et al., 2017). The Fermi-LAT analysis performed for the period in which the source was in the sub-luminous disk state (2008-2012) provides results consistent with those reported by Xing & Wang (2015) and Torres et al. (2017): XSS J1227-48538 is best described by a power-law with a cutoff at Ecut= 5.3 GeV (at 3.4) and a spectral index of = 2.
The Fermi-LAT analysis procedure performed on this source is similar to that described before for the other tMSP.
Similarly to PSR J1023+0038, the two spectral models considered for the CTAO simulations are the logparabola and the broken power-law. As input models for the CTAO simulations we considered the output from the Fermi-LAT study. For the broken power-law: Pf = (2.710.71) 10-12 photons cm-2 s-1 MeV-1, Index1 = -2.230.04, Index2 = -2.770.10, Energy break = (1.321.44) GeV. For the logparabola: Pf = (3.260.91) 10-11 photons cm-2 s-1 MeV-1, = 2.280.05, = 0.090.02, Energy break = (4.455.42) GeV. We performed batches of 100 simulations using only the CTAO-S site IRFs and 50, 100 and 200 h of observations. The simulated data were binned into 20 logarithmic energy bins in an energy range of 0.03-100 TeV.
The resulting spectrum of XSS J1227-48538 is shown in Figure 21 and, for simplicity, we reported only the results with the highest statistics. The complete results of the analysis for the broken power law are reported in Table 6 and the parameters of the unbinned analysis are:
Pf = 3.15 10-12 photons cm-2 s-1 MeV-1, = -2.23,
= -2.78, Energy break = 1.33 GeV. The source could be detected with CTAO-S at 8.14 with 50 h of observation.
As for PSR J1023+0038, we interpret the non convergence of the logparabola model as caused by the fact that the flux inferred from the fit falls marginally below the CTAO sensitivity curve for 200 hours.
Our simulations prove that the detection of the spectral component seen by Fermi-LAT from of close-by tMSPs when in a disc state could be possible with long exposures, provided that the emission has no cutoff at few GeVs. While this will likely not allow to catch fully a transition if lasting less than several days, here we demonstrate that once a transition has occurred CTAO could be able to detect such type of sources, identifying tMSPs as VHE emitters for the first time. While during the rotation-powered state MSP binaries will not be at the reach of CTAO, the known tMSPs have demonstrated an increase of their HE flux by a factor of 3-10 when transitioning from the radio to the disc states (see Torres et al. (2017)). The possibility to look-back in the CTAO Galactic Plane and other higher latitude pointings, where many MSP binaries are located, will be crucial for assessing any possibly related VHE flux increase or for complementing the X-rays and lower-energy coverage in case only upper limits are obtained. Also, if there are additional components (not considered here, such as magnetic reconnection of pulsar wind) it could be possible to detect changes in the VHE flux along transitions.
@@ 4.4 Flares in PWNe: the Crab Nebula
CTAO will cover a fundamental energy range to understand the origin of these flares: on the one hand, the low energy threshold will allow the sampling of the Fermi-LAT spectral shape at few tens of GeV of synchrotron nature, providing important clues on the acceleration and emission processes; on the other hand, the excellent sensitivity in the TeV regime will serve to explore the IC component that might arise at a detectable level from the electron population behind the MeV flares, off-scattering soft photon fields.
@@
To evaluate the capability of CTAO to detect Crab flares, we performed simulations of the SEDs both in flaring and steady (non-flaring) states of the nebula. We simulated flares of different spectral characteristics starting from a parent particle population, varying the physical properties of the environment. In particular, we simulated three types of flares: a very bright flare with a similar flux (at hundreds of MeV) to the one observed by Fermi-LAT in April 2011 (Buehler et al., 2012; Striani et al., 2013), which is the flare with the largest flux to date, and two dimmer flares corresponding to the first one re-scaled by a factor of 0.5 and 0.1. Thus, the corresponding flux enhancement at the simulated flares’ peaks (above 100 MeV of energy) ranges from 3 to 30 times the average flux of the nebula, as observed in tens of flares detected since 2007 (Huang et al., 2021). Since no spectral variability has been reported at HE, we assume the same spectral model for these dimmer flares.
The simulations of the nebula in flaring and steady states are performed for the CTAO-N with the methods and tools presented in previous works (Mestre et al., 2020, 2021). Mestre et al. (2021) pointed to the crucial role of the LSTs, best sensitive at the sub-TeV energy regime, in the prospects for the Crab flares’ detection (even in the early stages of CTAO-N operations). The latter, however, cast doubts on CTAO-S prospects due to the lack of LSTs in the Alpha configuration. . The electron population was simulated with a fixed index () of 2.5, to guarantee the detection in the TeV regime (e.g., from 1.25 TeV to 50 TeV) of the brightest model of flare in less than 10 h at 95% confidence level (see Fig. 6 of Mestre et al. 2021). The different flare models are computed for a magnetic field (B) in the acceleration region ranging from 100 G to 1 mG and compared (see section 2 of Mestre et al. 2021) to the steady nebula SED in both tens of GeV (e.g., from 20 GeV to 200 GeV) and TeV regimes. We obtained the simulations of the Crab nebula SED in steady state from Mestre et al. 2020. To compare the flaring and steady nebula simulations, we computed the mean total expected excess (e.g., counts from the source after background subtraction), both in flaring and steady state, in 21 bins of energy from 12 GeV to 200 TeV with observation times ranging from a few minutes to 500 h. We compared the excess distributions using the Pearson’s chi-squared test, corresponding the null hypothesis () to the steady state. Then, we consider the flare implies a detectable flux level if is rejected at significance (see Table 7; values in parentheses correspond to confidence).
@@
The simulations performed show that the different models of flare are best detected in the GeV regime and in particular in less than an hour at energies below 200 GeV, see Fig. 22. In the TeV regime, flares dimmer than April 2011 flare by a factor 0.5 (at hundreds of MeV) would be detected in less than 10 h for B G (see Table 7). As a reference, the gamma-ray flares from the Crab detected with Fermi-LAT showed varied duration scales from a few days to weeks and complex substructures, with sub-flares of a few hours or day-long duration (Huang et al., 2021). Some of the models considered imply an energy in TeV electrons larger than erg, being the synchrotron cooling time, and erg s-1, the luminosity of the nebula in gamma rays (Rudak & Dyks, 1998). However, note that the energy in electrons available in the nebula is not limited to the one computed above if particle re-acceleration takes place, which would introduce additional boosts in the electron population energy reservoir. The simulations performed, together with previous results reported (see Mestre et al., 2021), provide excellent prospects for detecting flares from the Crab Nebula with CTAO, especially for the LST subarray, featuring the best sensitivity at energies of a few tens of GeV.